OpenStudy (anonymous):

rite the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3)

OpenStudy (anonymous):

Since the new line is parallel to y=7x+2, the slope if the new line m=7. So for the new line slope intercept form: y=mx+c. y=7x+c To get c input the given point (1,-3) in the equation, - 3=7(1)+c, so c = - 10. Therefore slope intefcept form is y=7x-10.